OpenStudy (anonymous):

Which of the following is a quadratic function? A. y = x2 - 50 B. y = x3 – 40x C. y = x4 + 18x D. y = x – 81 Identify the quadratic function from the choices below whose parabola opens upward. A. y = 3x – 5x2 B. y = 4x2 – 6x – 2 C. y = 7x – 3x2 + 7 D. y = 100 – 4x2

OpenStudy (anonymous):

A. and B. assuming x2 = x^2

OpenStudy (anonymous):

Identify the quadratic function from the choices below whose parabola opens downward. A. y = -5x + x2 B. y = 2x2 – 4x C. y = x2 – 11x + 7 D. y = x – 2x2 + 9 Does the function f(x) = x2 - 2x + 2 have a maximum or minimum value? A. Maximum B. Minimum CAN YOU HELP ME WITH THESE TWO @shan4453

OpenStudy (anonymous):

D. and minimum